OpenStudy (pphalke):

two planes leave an airport at the same time, one flying east, the other flying west. the eastbound plane travels 110 mph slower. they are 3320 mi apart after 4 hr. find the speed of each plane. @janu16 its the same question just numbers changed

OpenStudy (janu16):

is it the same question but different numbers?

OpenStudy (pphalke):

yea they tricked me

OpenStudy (janu16):

so did oyou see the other answer?

OpenStudy (pphalke):

I saw it so far I did 4x+4(x+110=3320 is that right @Janu16

OpenStudy (janu16):

yes

OpenStudy (janu16):

4x + 4(x+110) = 3320 x=360 I defined x as the speed of the slower speed so the slower plane has speed of 360 mph I defined x+100 as the speed of the faster plane, so faster plane goes (360+110) or 470 mph check: what is 4(360) + 4(470)=3320 so the slower plane is 360mph fater plane is 470 mph

OpenStudy (janu16):

i got it. its up there^^

OpenStudy (anonymous):

470-360 is 110, not 100 .

OpenStudy (janu16):

i know i made a mistake cause i copied my last answer and forgot to change nubers

OpenStudy (pucusana):

could someone help me with a question?

OpenStudy (janu16):

she had a same question but different numbers on her last post so i forgot to chnage that numberr

OpenStudy (janu16):

@pucusana tag me and i will see

OpenStudy (anonymous):

\[{5 west + 5 east = 4000, east = west - 100} // Solve\]\[\{\text{east}=350,\text{west}=450\} \]

OpenStudy (pphalke):

@robtobey is ur answer right to my question

OpenStudy (janu16):

im right cause he has 4000 and remebr the number chnaged? to 3320?? @pphalke

OpenStudy (janu16):

@robtobey its not 4000 its 3320

OpenStudy (janu16):

@robtobey and its 110 not 100 you made error in the whole problem

OpenStudy (anonymous):

I believe that when the problem was first posted the distance apart was 4000 miles. Was the distance later changed to 3320 miles ?

OpenStudy (janu16):

yea cause these are 2 different problems

OpenStudy (janu16):

she got 2 problems with different numbers so she was confused by your answer

OpenStudy (anonymous):

Thank you for the response.

OpenStudy (janu16):

no problem